2. (a) Find the real root of the equation z3 + z + 10 = 0 given that one root is 1 − 2i.
Show that z = i is a root of the equation z4 + z3 + z − 1 = 0. Find the three
other roots.
1
Expert's answer
2020-05-26T18:13:19-0400

"z=1+2i" is also a root of equation. We obtain this by conjugating the equation. Dividing the polynomial "z^3+z+10" by "(z-1-2i)(z-1+2i)=z^2-2z+5"

we receive "z+2" . Thus, "z=-2" is the real root.


Now we consider equation "z^4+z^3+z-1=0" . We will check that "z_1=i" is a root of equation. Namely, we have: "z_1^4+z_1^3+z_1-1=1-i+i-1=0"

It is clear that "z_2=-i" is also a root of equation.

This can be obtained from the fact that "z_2=\bar{z}_1"

We shall divide the equation by "(z-i)(z+i)=z^2+1" .

Then one receives "z^2-1+z." It remains to solve "z^2+z-1=0" .

The latter quadratic equation has solutions (see e.g. https://en.wikipedia.org/wiki/Quadratic_equation for detalis)

"z_3 = \frac{-1+\sqrt{5}}2, \,\,z_4 = \frac{-1-\sqrt{5}}2" .


Thus, we have the roots "z_1=i,\quad z_2=-i,\quad z_3 = \frac{-1+\sqrt{5}}2,\quad z_4 = \frac{-1-\sqrt{5}}2,"